# TraveLER: A Multi-LMM Agent Framework for Video Question-Answering

### v1

@@ B.2 NExT-QA
Dataset. NExT-QA is a challenging dataset that tests causal action reasoning and temporal understanding. It contains 5,440 videos with an average length of 44s. Compared to earlier VideoQA benchmarks  (Papalampidi et al., 2023; Xue et al., 2017; Zeng et al., 2016; Zhu et al., 2015), NExT-QA requires going beyond simple recognition of objects and actions to answer the questions correctly. Each question requires selecting the best option out of 5 choices, often with very similar degrees of plausibility. Additionally, each question is categorized into either a Temporal, Causal, or Descriptive type. Temporal questions often ask what happens during, before, or after an event or action, while causal questions require advanced reasoning and inference about why an event or action occurs. Following the trend of works before us, we evaluate our method on the 5,000 questions in the NExT-QA validation set, which consist of 500 different videos.
@@ B.3 Perception Test
Dataset. In comparison with earlier VideoQA datasets (Maharaj et al., 2017; Zhu et al., 2015; Zeng et al., 2016) that focus on computational tasks such as classification, detection, or tracking, Perception Test is a dataset that focuses on skills such as memory, abstraction, physics, and semantics. Moreover, it is designed to test the transfer capabilities of different models and intended to be approached in a few-shot or zero-shot manner. The dataset consists of 11.6k real-world videos with an average length of 23 seconds, and 38K multiple choice QA questions.
@@
B.4 STAR
Dataset. STAR is a dataset that tests reasoning in real-world video situations. It consists of 22K video clips, with 60K situated reasoning questions, with 4 possible choices each. Questions are broadly divided into 4 main categories: interaction, sequence, prediction, and feasibility.
@@ Appendix C Additional Related Work
Modular Vision Frameworks. There has been a long history of work (Andreas et al., 2017; Herzig et al., 2018; Krishna et al., 2018; Baradel et al., 2018; Battaglia et al., 2018; Herzig et al., 2020; 2022; 2019; 2023; Bar et al., 2021; Avraham et al., 2022; Jerbi et al., 2020; Herzig et al., 2024; Mitra et al., 2023) that attempts to combine deep neural networks with modularity. Recently, works like VisProg (Gupta & Kembhavi, 2022), CodeVQA (Subramanian et al., 2023), RVP (Ge et al., 2023), and ViperGPT (Surís et al., 2023) have leveraged the improved coding capabilities of LMMs to generate code to compose different submodules together to answer visual questions. In addition, ProViQ (Choudhury et al., 2023) extends ViperGPT’s work in the video domain by adding more modules for VideoQA. Similarly, we leverage the strong power of LMMs in a modular approach. However, while these approaches have shown promising results, they are limited to single-shot planning when generating code, resulting in a fixed plan that cannot adapt. In contrast to these works, our approach has the advantage of being able to iteratively replan based on new information collected.
@@ Appendix D Qualitative Visualizations
We present further qualitative success and failure cases of our TraveLER framework. For each dataset, we display qualitative visualizations for 2 successes and 2 failure cases. For the success cases, we show expanded examples in Figures 5, 6, 7, and abridged versions in 8 that demonstrate the benefits of our question-answering approach compared to regular captioning. For failures, we also present abridged versions for each dataset in Figures 9, 10, and 11. Finally, we present some additional success and failure cases in 12. For the abridged versions, we display the video on top, and the traversal order using the numbered orange circles. In the row beneath, we display the frames in the order they are selected, and display corresponding Extractor question-answer output in yellow and captions in gray.

### v2

@@ B.2 NExT-QA
Dataset. NExT-QA is a challenging dataset that tests causal action reasoning and temporal understanding. It contains 5,440 videos with an average length of 44s. Compared to earlier VideoQA benchmarks  (Papalampidi et al., 2024; Xue et al., 2017; Zeng et al., 2017; Zhu et al., 2017), NExT-QA requires going beyond simple recognition of objects and actions to answer the questions correctly. Each question requires selecting the best option out of 5 choices, often with very similar degrees of plausibility. Additionally, each question is categorized into either a Temporal, Causal, or Descriptive type. Temporal questions often ask what happens during, before, or after an event or action, while causal questions require advanced reasoning and inference about why an event or action occurs. Following the trend of works before us, we evaluate our method on the 5,000 questions in the NExT-QA validation set, which consist of 500 different videos. This dataset is in English.
@@ B.3 Perception Test
Dataset. In comparison with earlier VideoQA datasets (Maharaj et al., 2017; Zhu et al., 2017; Zeng et al., 2017) that focus on computational tasks such as classification, detection, or tracking, Perception Test is a dataset that focuses on skills such as memory, abstraction, physics, and semantics. Moreover, it is designed to test the transfer capabilities of different models and intended to be approached in a few-shot or zero-shot manner. The dataset consists of 11.6k real-world videos with an average length of 23 seconds, and 38K multiple choice QA questions. This dataset is in English.
@@
B.4 EgoSchema
Dataset. EgoSchema is a dataset that tests long-form video reasoning and is intended to be answered in a few-shot or zero-shot manner. It introduces the idea of certificate lengths, which are the minimum number of seconds it takes to be able to answer the question correctly. It has 5k questions in the full set, and a 500 question subset is used as the validation set. We use the full test set for evaluation. This dataset is in English.
Inference Details. For EgoSchema, we use GPT-4 (gpt-4-1106-preview) as the LLM and LaViLa as the LMM. To prevent data leakage, we use a retrained version of LaViLa that does not use overlapping videos between Ego4D and EgoSchema. For window size in the Retriever, we take the selected frame, along with the frame 1 second before and 1-2 seconds after the selected frame.
B.5 STAR
Dataset. STAR is a dataset that tests reasoning in real-world video situations. It consists of 22K video clips, with 60K situated reasoning questions, with 4 possible choices each. Questions are broadly divided into 4 main categories: interaction, sequence, prediction, and feasibility. This dataset is in English.
@@ Appendix C Additional Related Work
Modular Vision Frameworks. There has been a long history of work (Andreas et al., 2016; Herzig et al., 2018; Krishna et al., 2018; Baradel et al., 2018; Battaglia et al., 2018; Herzig et al., 2020, 2022, 2019, 2023; Bar et al., 2021; Avraham et al., 2022; Jerbi et al., 2020; Herzig et al., 2024; Mitra et al., 2024) that attempts to combine deep neural networks with modularity. Recently, works like VisProg (Gupta and Kembhavi, 2023), CodeVQA (Subramanian et al., 2023), RVP (Ge et al., 2023), and ViperGPT (Surís et al., 2023) have leveraged the improved coding capabilities of LMMs to generate code to compose different submodules together to answer visual questions. In addition, ProViQ (Choudhury et al., 2023) extends ViperGPT’s work in the video domain by adding more modules for VideoQA. Similarly, we leverage the strong power of LMMs in a modular approach. However, while these approaches have shown promising results, they are limited to single-shot planning when generating code, resulting in a fixed plan that cannot adapt. In contrast to these works, our approach has the advantage of being able to iteratively replan based on new information collected.
@@ Appendix D Qualitative Visualizations
We present further qualitative success and failure cases of our TraveLER framework. For each dataset, we display qualitative visualizations for 2 successes and 2 failure cases. For the success cases, we show expanded visualizations in Figures Table 5, Table 6, Table 7, and abridged versions in 8 that demonstrate the benefits of our question-answering approach compared to regular captioning. For failures, we also present abridged versions for each dataset in Figures Table 9, 10, and 11. Finally, we present some additional success and failure cases in 12. For the visualizations, we display 3 iterations of the question answering process with 2 Extractor QA pairs each, compressing the output text by displaying the most important parts for brevity. For the abridged versions, we display the video on top, and the traversal order using the numbered orange circles. In the row beneath, we display the frames in the order they are selected, and display corresponding Extractor question-answer output in yellow and captions in gray.